# Bayesian post-correction of non-Markovian errors in bosonic lattice gravimetry

### 05 Mar 2026 v2

@@ Abstract
We study gravimetry with bosonic trapped atoms in the presence of random spatial inhomogeneity. The errors resulting from a random, shot-to-shot fluctuating spatial inhomogeneity are quantum non-Markovian. We show that in a system with $L>2$ modes (i.e., trapping sites), these errors can be post-corrected using a Bayesian inference. The post-correction is done via in situ measurements of the errors and refining the data-processing according to the measured error. We define an effective Fisher information $F_{\text{eff}}$ for such measurements with a Bayesian post-correction and show that the Cramer-Rao bound for the final precision is $\frac{1}{\sqrt{F_{\text{eff}}}}$. Exploring the scaling of the effective Fisher information with the number of atoms $N$, we show that it saturates to a constant when there are too many sources of error and too few modes. That is, with $\ell$ independent sources of error, we show that the effective Fisher information scales as $F_{\text{eff}}\sim\frac{N^{2}}{a+bN^{2}}$ for constants $a,b>0$ when the number of modes is small: $L<\ell+2$, even after maximization over the Hilbert space. With larger number of modes, $L\geq\ell+2$, we show that the effective Fisher information has a Heisenberg scaling $F_{\text{eff}}=\mathcal{O}(N^{2})$ when optimized over the Hilbert space. Finally, we study the density of the effective Fisher information in the Hilbert space and show that when $L\geq\ell+2$, almost any Haar random state has a Heisenberg scaling, i.e., $F_{\text{eff}}=\mathcal{O}(N^{2})$. Based on these results, we develop a loschmidt echo like experimental sequence for error mitigated gravimetry and gradiometry and discuss potential implementations. Finally, we argue that the effective Fisher information can be interpreted as the Fisher information corresponding to an equivalent non-Hertimitian evolution.
@@ I Introduction
The practical implementation of entangled homogeneous quantum sensors faces a significant challenge — with great sensitivity comes great fragility. That is, entangled states are highly sensitive to their environment, and therefore they are also very fragile and susceptible to multiple sources of noise and losses. Consequently, a practical quantum advantage in such sensors has largely remained elusive with the notable exception of LIGO [1].
@@
In particular, noise poses a significant hurdle to the scalability of quantum advantage in sensing. See Refs. [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] for recent experimental implementations of entanglement enhanced magnetometry with ultracold atoms. A practical quantum advantage requires a large number of atoms, typically, $N>10^{6}$. However, an entangled state with more than $10^{6}$ atoms is very susceptible to errors and therefore, retaining the quantum advantage is very challenging.
@@
Most of the work so far has been focused on Markovian errors, while non-Markovian errors are known to also have a significant impact on the scalability of entangled quantum sensors. Moreover, experimental work so far has been scarce and involves smaller number of qubits/atoms shedding no light on the impact of error correction on scalability. Finally, current works use a qubit-based architecture, whereas, many quantum sensing applications use bosons (e.g. Bose-Einstein Condensates, photons, etc).
@@ II Problem setup
We consider an atomic BEC with $N$ atoms in an optical lattice with $L$ sites, where each site can be occupied by multiple bosons (Fig. 2a,b). Hereafter, $\hat{a}_{i}^{\dagger}$ for $i=1,\cdots,L$ is the creation operator corresponding to site $i$. We assume that the Hamiltonian of the system is
@@ IV Correcting spatial inhomogeneity in gravimetry
Therefore, the $F_{\mathcal{N}}$ for almost any Haar state would be
@@
In Fig. 3a, b we show numerically computed $F_{\text{eff}}$ for a few examples. We use $\sigma_{\varphi}=\sigma_{i}=0.01$ fo all $i=1,\cdots,\ell$ and generate random errors where each $f_{ij}$ is uniformly random in [1,1][-1,1]. Fig. 3a clearly shows the asymptotic behavior of $F_{\text{eff}}$ — it saturates for $L<\ell+2$ and scales as $N^{2}$ for $L\geq\ell+2$. Fig. 3b shows how the effective Fisher information improves as we increase $L$, until it crosses $\ell+2$, after which it saturates. In the next section, we discuss a potential experimental implementation.
@@ V.1 Loschmidt echo sequence
The final state would be $U_{SP}\left|\psi_{0}\right$>. For a random choice of $J^{(k)}_{i},\Delta^{(k)}_{i}$ and $U^{(k)}_{i}$, we will almost certainly produce a state with $F_{\text{eff}}=\mathcal{O}(N^{2}/L^{2})$. Indeed, the typical value of the minimum eigenvalue of $F_{Q}$, $F_{\min}$ is,
@@
In Fig. 3c,d, we show that a random experimental pulse sequence indeed consistently produces a state with $F_{\text{eff}}$. In Fig. 3c, we show that the minimum eigenvalue of $F_{Q}$ after a random pulse sequence indeed follows Eq. (36) and therefore the effective Fisher information would scale as $N^{2}$. In Fig. 3d, we demonstrate the same, for larger values of $L$.
@@ VI Relation to non-Hermitian dynamics
One of the fruitful way to formulate measurement-feedback on a “leakage” space is non-Hermitian dynamics. In refs. [36], two-level non-Hermitian dynamics was implemented using a third “leakage” level. Population in this level was used to implement post-selected dynamics, which is effectively described by a non-Hermitian Hamiltonian. The post-correction developed here can be formulated as a non-Hermitian dynamics of an effective 2-level system described by the span of $\left|\psi\right$> and $H_{0}\left|\psi\right$>. The effective Fisher information $F_{\text{eff}}$ can be connected to the quantum Fisher information in non-Hermitian dynamics [37, 38]. Interestingly, better-than-SQL precision has been shown extensively in non-Hermitian systems [39, 40]. Using a non-Hermitian dynamics formulation for this work also allows us to expand its scope to errors originating from open system dynamics, which can be classically detected by monitoring photon emissions for instance.
@@ VII Conclusions
We have developed a technique for error mitigation in quantum sensing of gravity with multi-mode Bose-Einstein condensates. We address non-Markovian errors generated by random, fluctuating error Hamiltonians acting alongside gravity. We have shown that these errors can be mitigated using a Bayesian post-correction if the number of modes ($L$) and the number of independent sources of error ($\ell$) satisfy $L\geq\ell+2$. We also show that under this condition, the post-corrected precision has a Heisenberg scaling in the number of atoms — we can not only correct the errors, but we can also recover most of the quantum advantage using our technique. We developed a Loshmidt echo-based protocol to implement our technique experimentally.
@@ Acknowledgements
In this section, we provide a derivation of Eqs. 30 and 31 of the main text. Due to symmetry, we can use $i=1$ and $j=2$ in Eqs. 30 % 31 of the main text. Let

### 13 Apr 2026 v3

@@ Abstract
We study gravimetry with bosonic trapped atoms in the presence of random spatial inhomogeneity. The errors resulting from a random, shot-to-shot fluctuating spatial inhomogeneity are quantum non-Markovian. We show that in a system with $L>2$ modes (i.e., trapping sites), these errors can be post-corrected using a Bayesian inference. The post-correction is done via in situ measurements of the errors and refining the data-processing according to the measured error. We define an effective Fisher information $F_{\text{eff}}$ for such measurements with a Bayesian post-correction and show that the Cramer-Rao bound for the final precision is $\frac{1}{\sqrt{F_{\text{eff}}}}$. Exploring the scaling of the effective Fisher information with the number of atoms $N$, we show that it saturates to a constant when there are too many sources of error and too few modes. That is, with $\ell$ independent sources of error, we show that the effective Fisher information scales as $F_{\text{eff}}\sim\frac{N^{2}}{a+bN^{2}}$ for constants $a,b>0$ when the number of modes is small: $L<\ell+2$, even after maximization over the Hilbert space. With larger number of modes, $L\geq\ell+2$, we show that the effective Fisher information has a Heisenberg scaling $F_{\text{eff}}=\mathcal{O}(N^{2})$ when optimized over the Hilbert space. Finally, we study the density of the effective Fisher information in the Hilbert space and show that when $L\geq\ell+2$, almost any Haar random state has a Heisenberg scaling, i.e., $F_{\text{eff}}=\mathcal{O}(N^{2})$. Based on these results, we develop a Loschmidt echo like experimental sequence for error mitigated gravimetry and gradiometry and discuss potential implementations. Finally, we argue that the effective Fisher information can be interpreted as the Fisher information corresponding to an equivalent non-Hermitian evolution.
@@ I Introduction
With the rapid progress in quantum hardware technologies, achieving a quantum advantage in applications such as quantum sensing is more likely than ever. However, a definitive practical advantage has largely remained elusive, with a few notable exceptions including LIGO [1]. The major bottleneck hindering the scalability and field deployment of these sensors is their inherent susceptibility to environmental noise and operational errors. As hardware capabilities continue to mature, the development of robust error mitigation and correction protocols specifically tailored for sensing is gaining significant relevance.
@@
The conceptual tools for inference in quantum metrology remained largely frequentist, despite the tussle between Bayesian and frequentist inferences in general estimation theory being a few centuries old [8]. While Bayesian inference in quantum metrology is conceptually different from frequentist inference [9], the former gained relevance with mid-circuit measurements becoming a possibility. Since then, Bayesian inference has been used to develop adpative quantum sensing protocols [10, 11, 12] and experimental implementations [13, 14, 15].
Most of the work so far has been focused on Markovian errors, while non-Markovian errors are known to also have a significant impact on the scalability of entangled quantum sensors. Moreover, experimental work so far has been scarce and involves smaller number of qubits/atoms shedding no light on the impact of error correction on scalability. Finally, current works use a qubit-based architecture, whereas, many quantum sensing applications use bosons (e.g. Bose-Einstein Condensates, photons, etc). [16, 17, 18, 19, 20].
@@ II Problem setup
Trapped ultracold atoms have been a very successful platform for quantum sensing [21, 22]. We consider an atomic BEC with $N$ atoms in an optical lattice with $L$ sites, where each site can be occupied by multiple bosons (Fig. 2a,b). Hereafter, $\hat{a}_{i}^{\dagger}$ for $i=1,\cdots,L$ is the creation operator corresponding to site $i$. We assume that the Hamiltonian of the system is
@@ IV Correcting spatial inhomogeneity in gravimetry
Note that while these averages are calculated for the Haar measure, they only use upto the $4$-th moments and are valid for any 44-design. Therefore, the $F_{\mathcal{N}}$ for almost any state sampled from any 44-design would be
@@
In Fig. 3a, b we show numerically computed $F_{\text{eff}}$ for a few examples. We use $\sigma_{\varphi}=\sigma_{i}=0.01$ for all $i=1,\cdots,\ell$ and generate random errors where each $f_{ij}$ is uniformly random in [1,1][-1,1]. Fig. 3a clearly shows the asymptotic behavior of $F_{\text{eff}}$ — it saturates for $L<\ell+2$ and scales as $N^{2}$ for $L\geq\ell+2$. Fig. 3b shows how the effective Fisher information improves as we increase $L$, until it crosses $\ell+2$, after which it saturates. In the next section, we discuss a potential experimental implementation.
@@ V.1 Loschmidt echo sequence
The final state would be $U_{SP}\left|\psi_{0}\right$>. For a random choice of $J^{(k)}_{i},\Delta^{(k)}_{i}$ and $U^{(k)}_{i}$, we will almost certainly produce a state with $F_{\text{eff}}=\mathcal{O}(N^{2}/L^{2})$ [23]. Indeed, the typical value of the minimum eigenvalue of $F_{Q}$, $F_{\min}$ is,
@@
In Fig. 3c,d, we show that a random experimental pulse sequence indeed consistently produces a state with $F_{\text{eff}}$. In Fig. 3c, we show that the minimum eigenvalue of $F_{Q}$ after a random pulse sequence indeed follows Eq. (36) and therefore the effective Fisher information would scale as $N^{2}$. In Fig. 3d, we demonstrate the same, for larger values of $L$. In ref. [23], we have provided a mathematical proof that the states produced by random evolutions under Hamiltonians of the form Eq. (34) approach a 44-design and therefore, states produced under this evolution almost certainly have a QFI given by Eq. (32). Accordingly, for these states, $F_{\text{eff}}=\mathcal{O}(N^{2}/L^{2})$.
@@ VI Relation to non-Hermitian dynamics
One of the fruitful ways to formulate measurement-feedback on a “leakage” space is non-Hermitian dynamics. In refs. [24], two-level non-Hermitian dynamics was implemented using a third “leakage” level. Population in this level was used to implement post-selected dynamics, which is effectively described by a non-Hermitian Hamiltonian. The post-correction developed here can be formulated as a non-Hermitian dynamics of an effective 2-level system described by the span of $\left|\psi\right$> and $H_{0}\left|\psi\right$>. The effective Fisher information $F_{\text{eff}}$ can be connected to the quantum Fisher information in non-Hermitian dynamics [25, 26]. Interestingly, better-than-SQL precision has been shown extensively in non-Hermitian systems [27, 28]. Using a non-Hermitian dynamics formulation also allows us to expand the scope of this work to errors originating from open system dynamics, which can be classically detected by monitoring photon emissions for instance.
@@ VII Conclusions
We have developed a technique for error mitigation in quantum sensing of gravity with multi-mode Bose-Einstein condensates. We address non-Markovian errors generated by random, fluctuating error Hamiltonians acting alongside gravity. We have shown that these errors can be mitigated using a Bayesian post-correction if the number of modes ($L$) and the number of independent sources of error ($\ell$) satisfy $L\geq\ell+2$. We also show that under this condition, the post-corrected precision has a Heisenberg scaling in the number of atoms — we can not only correct the errors, but we can also recover most of the quantum advantage using our technique. We developed a Loschmidt echo-based protocol to implement our technique experimentally.
@@ Acknowledgements
In this section, we provide a derivation of Eqs. 30 and 31 of the main text. Due to symmetry, we can use $i=1$ and $j=2$ in Eqs. 30 & 31 of the main text. Let